ABSOLUTE LINE INTENSITIES OF H_{2}CO IN THE 3.5 AND 5.7-μm REGIONS

\underline{F. KWABIA TCHANA}, N. LACOME, Université Pierre et Marie Curie-Paris 6; CNRS; Laboratoire de Dynamique, Interactions et Réactivité, UMR 7075, Case Courrier 49, 4 Place Jussieu, 75252 Paris Cedex 05, France; A. PERRIN, Laboratoire Inter Universitaire des Systèmes Atmosphériques (LISA), CNRS, Universités Paris 12 et 7, 61 Avenue du Général de Gaulle, 94010 Créteil Cedex, France.

Formaldehyde (H_{2}CO) is detected in the earth’s troposphere by infrared techniques in the 3.5 and 5.7-μm regions. Recent measurements produced line positions and relative line intensities^{a} and also IR⇔UV line intensity intercomparisons ^{b}. For atmospheric retrievals absolute intensities and line broadening parameters are needed. For that, new Fourier transform spectra were recorded at high resolution (0.0035 cm^{−1}) at LADIR in the 1600-3000 cm^{−1} spectral region. Low pressures (up to 0.5 torr) of H_{2}CO were generated by warming paraformaldehyde. An appropriate monitoring of the heating temperature (≈ 40^{∘} C) avoided any polymerization and allowed to obtained a stable pressure of pure H_{2}CO (98 ± 1%). In this way accurate line positions and absolute intensities of H_{2}CO were measured and theoretical modelled in the 3.5 and 5.7-μm regions leading also to an intercomparison of intensities between the two spectral regions. The determination of self- and N_{2}-broadening coefficients is in progress.

aA. Perrin, F. Keller, and J.-M. Flaud, J. Mol. Spectrosc., 221, (2003) 192. A. Perrin, A. Valentin, and L. Daumont, J. Mol. Struct., 780-781, (2006) 28.

bA. Gratien, B. Picquet-Varrault, J. Orphal, J.-F. Doussin, and J.-M. Flaud, J. Geophys. Res. D, in press, (2007).